50 ml of 1M oxalic acid is shaken with 0.5 g of wood charcoal. The final concentration of solution after adsorption is 0.5 M .Calculate the amount of oxalic acid adsorbed per gram of charcoal?

Grade:12th pass

6 Answers

bharat bajaj IIT Delhi
10 years ago
Total moles of oxalic acid = 50 mmol
Final concentration = 0.5 M
Final no. of moles = 0.5/50 ml = 10 mmol
amount of oxalic acid adsorbed per gram = 40 mmol/0.5 = 80 mmol/g = 80 X 126 / 1000 = 10.08 gram of oxalic acid / gram of charcoal
Thanks
Bharat Bajaj
IIT Delhi
askiitians faculty
Vinu nayak
6 years ago
Initial concentration =1MFinal concentration = 0.5 MVolume = 0.05 LInitial no of moles = 1/20Initial mass of oxalic acid = 6.3 gFinal mass of oxalic acid = 3.15gAmount of oxalic acid adsorbed per gram of charcoal =3.15/0.5 =6.3gVinu Thanks
Vatsala jandotia
6 years ago
1M oxalic acid solutions mean; 1mole of oxalic acid presents in 1000mL solution 50mL of 1M solutions will contain oxalic acid =126/1000×50=6.6g
SHEEBA
6 years ago
Initial amount of oxalic acid present in 50 ml=126/1000×50=6.3g.Amount of oxalic acid in 50ml after adsorption=3.15gAmount of oxalic acid adsorbed per gram of charcoal =3.15×2=6.3g
Sanskar Goyal
5 years ago
Given, Volume of Oxalic Acid = 50ml = 0.05L
Initial Concentration- 1M
Final Concentration- 0.5M
So, Initial Moles- (0.05*1)= 0.05moles
Final Moles- (0.05*0.5)= 0.025moles
Change in Moles= 0.05-0.025 = 0.025moles
 
 
Now, Amount adsorbed=
\frac{(Change~in~Moles)*(Molecular~Mass)}{Grams~of~Carbon}So, (0.025*90.03)/0.5  = 4.5015 grams
Thanks,
Sanskar Goyal
Kushagra Madhukar
3 years ago
Dear student,
Please find the attached solution to the problem.
 
Initial moles of oxalic acid = 1M * 0.05L = 0.05 moles
Final concentration = 0.5M * 0.05L = 0.025 moles
Moles of oxalic acid adsorbed = 0.05 – 0.025 = 0.025 moles
Molar mass of oxalic acid = 90g
Mass of oxalic acid adsorbed = 0.025 * 90 = 2.25g
Hence oxalic acid adsorbed per gram of charcoal = 2.25/0.5 = 4.5
 
Hope it helps.
Thanks and regards,
Kushagra